OpenStudy (anonymous):

additive inverse and multiplicative inverse of 2.5

jimthompson5910 (jim_thompson5910):

additive inverse solve for x: x+2.5 = 0

jimthompson5910 (jim_thompson5910):

multiplicative inverse solve for y 2.5*y = 1

OpenStudy (anonymous):

what would y be?

jimthompson5910 (jim_thompson5910):

you need to isolate y in 2.5*y = 1 to find out

jimthompson5910 (jim_thompson5910):

to do that, you divide both sides by 2.5

OpenStudy (anonymous):

so y would be 1/2.5?

jimthompson5910 (jim_thompson5910):

evaluate that and you will get your answer

OpenStudy (anonymous):

ok thank you

jimthompson5910 (jim_thompson5910):

you're welcome